ir a medicine has a dosage of 20ml how many doses are in 4 gallon bottle
Answer #1

1 gallon = 3785 ml

4 gallon = 3785 *4 ml

By unitery method ,

20 ml = 1 dose

1 ml = 1/20 dose

3785 *4 ml =( 1 *4* 3785 ) / 20 dose

= 757 dose

Ans = 757 dose